OpenStudy (anonymous):

A missing fraction on a number line is located exactly halfway between 3/6 and5/6.what is the missing fraction?

OpenStudy (anonymous):

how do you think this can b solved?

OpenStudy (jdoe0001):

\(\large \begin{array}{lllllll} \frac{1}{6}&\frac{2}{6}&\frac{3}{6}&\frac{\square }{\square }?&\frac{5}{6}&\frac{6}{6}& \end{array} \)